# GTPO: Trajectory-Based Policy Optimization in Large Language Models

### v3

@@ 4.2 Policy Collapse
The GRPO gradient, as discussed in Eq. (2), highlights a dependency on both the advantage and the probability scores assigned by the policy $\pi_{\theta}$. To gain a deeper understanding of how the advantage influences the learning dynamics, we analyze the average Shannon entropy of the output distribution $\pi_{\theta}$ over completions $i$ and tokens $t$ during training. The Shannon entropy is defined as $H(\mathbf{p})=-\sum_{j=1}^{n}p_{j}\ln p_{j},$ where $\mathbf{p}$ is a probability vector. For a token position $t$ in a completion $o_{i}$, we consider $\mathbf{p}=\pi_{\theta}(o_{i,t}|s_{i,t})$, and we define $\langle H\rangle_{i}$ as the average entropy across in the completion $o_{i}$.
At a generation step, let us consider a low entropy (e.g., $\langle H\rangle_{i}\ll\ln 2$)  [^2] A completion with average entropy near $\ln 2$ suggests balanced uncertainty between two choices for each output token. , indicating that the model is highly confident, i.e., it assigns almost all probability mass to a single token index $j$, with $\pi_{\theta}^{j}\approx 1$ and $\pi_{\theta}^{k}\approx\epsilon\ll 1$, where $k\neq j$. If this token leads to a negative outcome ($\hat{\mathcal{A}}_{i}<0$), GRPO penalizes it sharply, since the gradient of the GRPO loss at index $\tau$ for completion $i$ becomes:
@@
$\nabla_{\theta}{\mathcal{J}}^{(i,\tau)}_{\text{GRPO}}(\theta)\approx\begin{cases}-\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}\cdot\nabla_{\theta}f_{\theta}^{j}\cdot(1-\epsilon)&\text{for }j\\ \;\;\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}\cdot\nabla_{\theta}f_{\theta}^{k}\cdot\epsilon&\text{for }k\neq j\end{cases}$ (3)
Note that the KL term has been omitted here, its contribution will be addressed next.
This results in a negative update for the selected token $j$, and small positive updates for all other tokens $k\neq j$, even if they are potentially syntactically or semantically incorrect. While each of these positive updates is individually small, they can accumulate over time, gradually increasing the probability of initially implausible tokens.
@@
The effect is further amplified by GRPO’s zero-mean constraint: when most completions receive positive rewards, the few with negative rewards must carry disproportionately large negative advantages to maintain balance. These penalties can suppress correct predictions and unintentionally amplify the likelihood of undesirable alternatives, raising entropy and introducing instability. Over time, this harms output quality and increases the risk of introducing structural and semantic errors that propagate through subsequent training steps. As the model drifts away from desirable behaviors, the reward signal becomes less meaningful.
@@
Figure 1: Training behavior of LLaMA-8B on GSM8K using two GRPO runs with $\beta=0.04$ and $\beta=10^{-6}$, both with $G=8$. The top plot shows the formatting rewards for both runs, and the bottom plots show the average entropy and KL divergence in each run.
This phenomenon is illustrated in the top plot of Figure 1, which shows the average formatting reward during the training of LLaMA-8B on GSM8K, using GRPO with $\beta=0.04$ (as originally proposed in (Shao et al. 2024)) and $\beta=10^{-6}$. At the beginning of training, the formatting reward increases more rapidly with $\beta=10^{-6}$, while it remains relatively flat with $\beta=0.04$. However, in both cases, the reward begins to drop below 9, after 580 steps for $\beta=10^{-6}$ and around step 750 for $\beta=0.04$.
@@
KL Reacts Late to Avoid Collapse. Middle and bottom plots of Figure 1 show the average entropy $\langle H\rangle_{i}$ and average KL divergence, computed on the generated completions, for the two previous training runs of LLaMA 8B. Both KL divergence curves remain stable up to approximately step 1200, well after the degradation points, when the formatting rewards have already collapsed and entropy levels are high. Only around the KL peak, the reward curves show signs of stabilization. In contrast, the entropy curves begin to rise sharply as the formatting rewards start to decline, and continue to increase as the rewards deteriorate further.
This pattern suggests a key insight: KL divergence acts as a delayed corrective signal, rising only after collapse. In contrast, entropy tracks the policy collapse in real time, increasing as the policy loses structure. Further analyses of the entropy behaviour are provided in the appendix.
@@ 5 Method
This section introduces GTPO, which addresses the previous issues by (i) mitigating gradient conflict on shared tokens, and (ii) preventing policy collapse during training by regularizing the model behavior through the entropy.
@@ 5.1 Conflict-Aware Gradient Correction
As discussed in Section 4.1, The model can encounter gradient conflict issues, where shared tokens receive both positive and negative updates depending on the advantage assigned to each completion. To correct this issue, we propose in the following a methodology aimed at selectively masking gradient contributions for tokens involved in conflicting updates. To this end, we first identify conflict tokens, as tokens that appear in the same position, either from the beginning or the end, across completions with both positive and negative advantages. As discussed in the context of GRPO’s issues part, these tokens often receive conflicting gradient updates during training. Then, we mitigate such conflicts by correcting their gradient updates accordingly.
Conflict tokens definitions. Let $\{o_{1},\dots,o_{|G|}\}$ be a group of completions. Each completion $o_{i}$ contains a sequence of tokens $\{o_{i,1},\dots,o_{i,|o_{i}|}\}$, and is associated with an advantage value $A^{(i)}\in\mathbb{R}$. We define $G^{-}$ and $G^{+}$ as the sets of completions with $\mathcal{A}<0$ and $\mathcal{A}>0$, respectively.
Left-to-right alignment. A token $v\in\mathcal{V}$ is a forward conflict token at position $p$ if it appears at the same position in at least one completion with positive advantage and in at least one with negative advantage:
@@
Gradient Reweighting with conflict masks. Based on the definitions above, we construct binary masks for tokens across completions to target positions potentially affected by gradient conflicts and thus correct their updates.
We first define a forward mask $\mathcal{M}^{\mathrm{fw}}_{i}\in\{0,1\}^{|o_{i}|}$ by scanning $o_{i}$ from left to right, setting 1 over the first contiguous span of forward conflict tokens and 0 elsewhere. Analogously, the backward mask $\mathcal{M}^{\mathrm{bw}}_{i}$ is obtained by scanning from right to left, marking the first contiguous span of backward conflict tokens. A final mask $\mathcal{M}_{i}$ is then defined as: $\mathcal{M}_{i}=\mathcal{M}^{\mathrm{fw}}_{i}\lor\mathcal{M}^{\mathrm{bw}}_{i},$ highlighting only the initial and final conflict regions in each completion $o_{i}$.
After computing each $\mathcal{M}_{i}$ in the group of completions, we correct inconsistent gradient updates over the selected conflict tokens, while leaving other token updates unchanged, thought the following preliminary token-level loss:
$\mathcal{J}^{*}=\frac{1}{G}\sum_{i=1}^{G}\frac{\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\lambda_{i,t}$ (4)
@@
Intuitively, the mask disables negative gradients on conflict tokens, and instead reinforces them only if they appear in positively rewarded completions. Note that the total signal magnitude is preserved across the group: since $\sum_{i\in G^{+}}|\mathcal{A}_{i}|=\sum_{i\in G^{-}}|\mathcal{A}_{i}|$, doubling the signal for positive completions compensates for the removal of negative updates, maintaining training stability while preventing semantic drift. Additional details of the weighting scheme are provided in appendix, while its benefits are evaluated in the ablation studies (Section 6.2).
Importantly, the final mask $\mathcal{M}_{i}$ targets only the initial and final contiguous conflict tokens to preserve the semantic structure of completions. In fact, masking isolated conflict tokens in the middle could harm stability and learning, as their meaning often depends on surrounding context, and altering their gradients may be counterproductive. In contrast, the outer spans typically correspond to formatting tags, such as <reasoning> or </answer>, which are the primary source of conflict in GRPO, as discussed in Section 4.1. Focusing the correction on these regions protects structural tokens without interfering with the central part of the completion, where meaningful differences in trajectories emerge.
5.2 Entropy-Based Policy Regularization
@@
Entropy Regulatization. Inspired by PPO (Andrychowicz et al. 2021; Huang et al. 2022), we add a regularization term based on the average tokens entropy of each completion, $\langle H\rangle_{i}$, where $\gamma$ balance the importance of this term in the final loss, as shown below. Note that, based on the internal characteristics of GRPO to implicitly increase entropy over time, we decided to minimize the term. This acts as a way to reduce the model entropy over time.
The combination of these entropy-based strategies and the token-level loss defines the final GTPO objective, as:
$\mathcal{J}_{\text{GTPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{\delta_{i}\cdot\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\lambda_{i,t}\;-\;\gamma\cdot\langle H\rangle_{i}$ (7)
The proposed GTPO loss does not require an additional reference model, unlike the KL divergence term in GRPO, thereby reducing the memory footprint during training. The benefits of each component, and hyperparameter of the loss are discussed and demonstrated, also through ablation studies in the following experimental section.
@@ 6 Experiments
Experimental Setup. We conducted experiments to assess the training stability and generalization performance of GTPO. All experiments were performed on LLaMA-8B (Patterson et al. 2022) and Qwen 2.5 (3B) (Yang et al. 2025a), which have trained using the training splits of GSM8K (Hendrycks et al. 2021a) and MATH (Hendrycks et al. 2021b), and evaluated after training on the corresponding test splits, and also the AIME2024 dataset (AIME 2024).
Figure 3: Out-of-distribution evaluation on AIME2024, with pass@k (%) over $k$, on models trained on MATH and GSM8K.
For comparison, all models were also trained using SFT and GRPO (with both $\beta=0$ and $\beta=10^{-6}$ to assess the impact of the KL term). [^3] Additional analysis of different $\beta$ values is provided in the appendix; $\beta=10^{-6}$ was selected as it yields the best results. To further explore hyperparameter group-relative optimizations, both GRPO and GTPO were evaluated with two generation sizes: $G=8$ and $G=12$. For the entropy-based completion mask used in GTPO, we compute $\langle H\rangle_{\text{ini}}$ by evaluating the original LLM entropy on the first 100 samples of the training set.
All training was performed using a learning rate of $10^{-6}$, with the temperature set to 1.0 during the test phase. Further details of the experimental setting are in the Appendix. All experiments were conducted on 2 NVIDIA A100 GPUs.
@@ 6.1 Performance Evaluation
Training Dynamics of GTPO. In Figure 2-top (the first two rows), we compare the training stability and performance of GTPO against GRPO. Formatting and accuracy rewards are reported as percentages, where the maximum value (100%) corresponds to a reward of 10 in both. On the GSM8K dataset with LLaMA, GTPO consistently outperforms GRPO across all training steps in both accuracy and formatting metrics. On the more challenging MATH dataset, GRPO (with $G=12$ and $\beta=0$) initially achieves slightly better accuracy around the midpoint of training. However, its performance drops sharply in the second half of training due to policy collapse, affecting both accuracy and formatting. In contrast, GTPO continues to improve steadily throughout training, avoiding collapse and maintaining stable performance. For Qwen 2.5, on both GSM8K and MATH, GTPO achieves comparable or improved accuracy relative to GRPO, with only a slight decrease in formatting performance (still above 97% in all runs). Note that, consistent with the analysis in Section 4.2, GRPO training curves for Qwen 2.5 are not affected by a policy collapse, as it exhibits high-entropy behavior. Overall, GTPO demonstrates more stable and reliable training compared to GRPO across models and datasets.
In-distribution Evaluation. The trained LLMs were evaluated on the test sets of GSM8K and MATH using the pass@k (Chen et al. 2021) and maj@k (Wang et al. 2023) metrics. The former measures whether at least one of the top-$k$ completions yields a correct answer, while the latter assesses correctness via majority voting over the top-$k$ completions. Note that, to ensure a fair comparison, we evaluated GRPO on LLaMA using the model checkpoint corresponding to the training step with the highest accuracy reward, rather than the one obtained after policy collapse.
@@
Figure 2-bottom (the last two rows) shows that GTPO consistently outperforms GRPO in almost all settings for both pass@k and maj@k, as $k$ varies from 1 to 32. This indicates that models trained with GTPO exhibit stronger self-consistency when answering questions (higher maj@k) and better coverage of the correct answer across multiple completions (higher pass@k). We also include testing comparisons with SFT, where GTPO consistently outperforms SFT in maj@k across all models and datasets, and achieves higher average performance in pass@k. Specifically, SFT surpasses GTPO only in pass@k for $k>5$ on MATH with LLaMA, but not in terms of correctness with maj@k. Interestingly, in GTPO, larger values of $G$ lead always to better performance, which is not the same for GRPO.
Figure 4: Analysis of training dynamics under different settings of entropy regularization $\gamma$: average group completion entropy (left), accuracy (center), and formatting (right). Note that $\gamma=0.1$, No $\delta_{i}$’ denotes the setting in which the entropy-based filtering term is disabled.
@@
Figure 5: Accuracy and formatting performance of LLaMA trained on GSM8K. For GRPO, results are shown under varying KL-$\beta$ values. For GTPO, “No $\langle H\rangle_{i}$ - No $\delta_{i}$" denotes the setting in which only the Conflict-aware Gradient Correction is applied.
Out-of-distribution Evaluation. We also evaluate the trained models on a different test set (AIME2024), as shown in Figure 3, reporting pass@k scores with $k$ extended up to 64 to account for the increased difficulty of the task. For convenience, for each dataset, we select the GRPO and GTPO variants with the generation size $G$ that yielded the best performance on the in-distribution tests.
The results show that GTPO consistently outperforms both SFT and GRPO, particularly at higher values of $k$ on the MATH dataset, where the increased complexity encourages broader exploration of reasoning paths. In contrast, GRPO does not consistently benefit from this, which limits its performance gains. Interestingly, both GTPO and GRPO demonstrate stronger out-of-distribution generalization compared to SFT, suggesting a higher degree of overfitting to the in-distribution data in the latter.

### v4

@@ 4.2 Policy Collapse
The GRPO gradient, Eq. (3), highlights a dependency on both the advantage and the probability scores assigned by the policy $\pi_{\theta}$. To gain a deeper understanding of how the advantage influences the learning dynamics, we analyze the average Shannon entropy of the output distribution $\pi_{\theta}$ over completions $i$ and tokens $t$ during training. The Shannon entropy is defined as $H(\mathbf{p})=-\sum_{j=1}^{n}p_{j}\ln p_{j},$ where $\mathbf{p}$ is a probability vector. For a token position $t$ in a completion $o_{i}$, we consider $\mathbf{p}=\pi_{\theta}(o_{i,t}|s_{i,t})$, and we define $\langle H\rangle_{i}$ as the average entropy across in the completion $o_{i}$. At a generation step, let us consider a low entropy (e.g., $\langle H\rangle_{i}\ll\ln 2$)  [^2] An average entropy close to $\ln 2$ indicates that the model is choosing between two equiprobable tokens at each step. , indicating that the model is highly confident, i.e., it assigns almost all probability mass to a single token index $j$, with $\pi_{\theta}^{j}\approx 1$ and $\pi_{\theta}^{k}\approx\epsilon\ll 1$, where $k\neq j$. If this token leads to a negative outcome ($\hat{\mathcal{A}}_{i}<0$), GRPO penalizes it sharply, since the gradient of the GRPO loss at index $\tau$ for completion $i$ becomes:
@@
$\nabla_{\theta}{\mathcal{J}}^{(i,\tau)}_{\text{GRPO}}(\theta)\approx\begin{cases}-\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}r_{j,i}(\theta)\nabla_{\theta}f_{\theta}^{j}\cdot(1-\epsilon)&\text{for }j\\ \;\;\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}r_{k,i}(\theta)\nabla_{\theta}f_{\theta}^{k}\cdot\epsilon&\text{for }k\neq j\end{cases}$ (7)
@@
Note that the KL term has been omitted here, its contribution will be addressed next. Both $r_{k,i}(\theta)$ and $r_{j,i}(\theta)$ are equal to $1$ at the first iteration and remain approximately $1$ in the subsequent ones. This results in a negative update for the selected token $j$, and small positive updates for all other tokens $k\neq j$, even if they are potentially syntactically or semantically incorrect. While each of these positive updates is individually small, they can accumulate over time, gradually increasing the probability of initially implausible tokens. The effect is further amplified by GRPO’s zero-mean constraint (the average of the advantages of the group $\bar{A}=0$): when most completions receive positive rewards, the few with negative rewards must carry disproportionately large negative advantages to maintain balance. These penalties can suppress correct predictions and unintentionally amplify the likelihood of undesirable alternatives, raising entropy and introducing instability. Over time, this harms output quality and increases the risk of introducing structural and semantic errors that propagate through subsequent training steps. As the model drifts away from desirable behaviors, the reward signal becomes less meaningful. This phenomenon is illustrated in the top plot of Figure 1, which shows the average formatting reward during the training of LLaMA 8B on GSM8K, using GRPO with $\beta=0.04$ (as originally proposed in GRPO) and $\beta=10^{-6}$. At the beginning of training, the formatting reward increases more rapidly with $\beta=10^{-6}$, while it remains relatively flat with $\beta=0.04$. However, in both cases, the reward begins to drop below 9, after 580 steps for $\beta=10^{-6}$ and around step 750 for $\beta=0.04$.
@@
Figure 1: GRPO training of LLaMA 8B on GSM8K $G=8$ comparing $\beta=0.04$ and $\beta=10^{-6}$. Plots show formatting reward (top), average entropy (middle), and KL divergence (bottom).
@@
KL Reacts Late to Avoid Collapse. Middle and bottom plots of Figure 1 show the average entropy $\langle H\rangle_{i}$ and average KL divergence, computed on the generated completions, for the two previous training runs of LLaMA 8B. Both KL divergence curves remain stable up to approximately step 1200, well after the degradation points, when the formatting rewards have already collapsed and entropy levels are high. Only around the KL peak, the reward curves show signs of stabilization. In contrast, the entropy curves begin to rise sharply as the formatting rewards start to decline, and continue to increase as the rewards deteriorate further. This suggests that the KL divergence acts as a delayed corrective signal, rising only after collapse. In contrast, entropy tracks the policy collapse in real time, increasing as the policy loses structure.
@@ 5 Method
This section presents GTPO, which mitigates gradient conflicts and prevents policy collapse.
@@ 5.1 Conflict-Aware Gradient Correction
As discussed in Section 4.1, The model can encounter gradient conflict issues, where shared tokens receive both positive and negative updates depending on the advantage assigned to each completion. Thus, we propose a methodology aimed at selectively masking gradient contributions for tokens involved in conflicting updates. To this end, we first identify conflict tokens, as tokens that appear in the same position, either from the beginning or the end, across completions with both positive and negative advantages. As discussed in the context of GRPO’s issues part, these tokens often receive conflicting gradient updates during training. Then, we mitigate such conflicts by correcting their gradient updates accordingly.
Conflict tokens definitions. Let $\{o_{i}\}_{i=1}^{G}$ be a group of completions. Each completion $o_{i}$ contains a sequence of tokens $\{o_{i,t}\}_{t=1}^{|o_{i}|}$, and is associated with an advantage value $A^{(i)}\in\mathbb{R}$. We define $G^{-}$ and $G^{+}$ as the sets of completions with $\mathcal{A}<0$ and $\mathcal{A}>0$, respectively.
Left-to-right alignment. A token $v\in\mathcal{V}$ is a forward conflict token at position $p$ if it appears at the same position in at least one completion with positive advantage and in at least one with negative:
@@
Gradient Reweighting with conflict masks. We construct binary masks for tokens across completions to target positions potentially affected by gradient conflicts and thus correct their updates. We first define a forward mask $\mathcal{M}^{\mathrm{fw}}_{i}\in\{0,1\}^{|o_{i}|}$ by scanning $o_{i}$ from left to right, setting 1 over the first contiguous span of forward conflict tokens and 0 elsewhere. The backward mask $\mathcal{M}^{\mathrm{bw}}_{i}$ is obtained by scanning from right to left, marking the first contiguous span of backward conflict tokens. A final mask $\mathcal{M}_{i}$ is then defined as: $\mathcal{M}_{i}=\mathcal{M}^{\mathrm{fw}}_{i}\lor\mathcal{M}^{\mathrm{bw}}_{i},$ highlighting only the initial and final conflict regions in each completion $o_{i}$. After computing each $\mathcal{M}_{i}$, we correct inconsistent gradient updates over the selected conflict tokens, while leaving other token updates unchanged, thought the following preliminary token-level loss:
$\mathcal{J}^{*}=\mathbb{E}_{q,\{o_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}r_{i,t}\lambda_{i,t}\right]$ (8)
@@
Intuitively, the mask disables negative gradients on conflict tokens, and instead reinforces them only if they appear in positively rewarded completions. Note that the total signal magnitude is preserved across the group: since $\sum_{i\in G^{+}}|\mathcal{A}_{i}|=\sum_{i\in G^{-}}|\mathcal{A}_{i}|$, doubling the signal for positive completions compensates for the removal of negative updates, maintaining training stability while preventing semantic drift. Additional details and derivations of the weighting scheme are provided in appendix 8.2, while its benefits are evaluated in the ablation studies (Section 6.2). Importantly, the final mask $\mathcal{M}_{i}$ targets only the initial and final contiguous conflict tokens to preserve the semantic structure of completions. In fact, masking isolated conflict tokens in the middle could harm stability and learning, as their meaning often depends on surrounding context, and altering their gradients may be counterproductive. In contrast, the outer spans typically correspond to formatting tags, such as <reasoning> or </answer>, which are the primary source of conflict in GRPO, as discussed in Section 4.1. Focusing the correction on these regions protects structural tokens without interfering with the central part of the completion, where meaningful differences in trajectories emerge.
5.2 Entropy Control
@@
Entropy Regularization. Inspired by PPO entropy-on-ppo; ppo37implementations, we add a regularization term based on the average tokens entropy of each completion, $\langle H\rangle_{i}$, where $\gamma$ balance the importance of this term in the final loss, as shown below. Note that, based on the internal characteristics of GRPO to implicitly increase entropy over time, we decided to minimize the term. This acts as a way to reduce the model entropy over time. The combination of these entropy-based strategies and the token-level loss defines the final GTPO objective, as:
$\mathcal{J}_{\text{GTPO}}=\mathbb{E}_{q,\{o_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\delta_{i}\cdot\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}r_{i,t}\lambda_{i,t}\;-\;\gamma\cdot\langle H\rangle_{i}\right]$ (11)
In contrast to vanilla-GRPO GRPO, shown in Equation 1, the proposed GTPO Objective does not rely on KL divergence, so the reference model is not required during training, making the process more lightweight and faster. Others hu2025open; liu2025understanding tried to remove the KL term from GRPO objective, but our tests (see Section 6) show it is sometimes necessary to get a more reliable and stable GRPO training. The benefits of each component, and hyperparameter of GTPO Objective are discussed and demonstrated, also through ablation studies in the following experimental section.
@@ 6 Experiments
[…]
AIME 2024 AIME 2025 AMC 2023
Pass@k (%) Pass@k (%) Pass@k (%) Maj@k (%)
Model Dataset Method 1 8 16 32 64 1 8 16 32 64 1 8 16 32 64 4 8 16 32 64
LLAMA
GSM8K SFT 0.0 0.0 0.0 13.3 20.0 0.0 0.0 0.0 0.0 6.7 12.5 37.5 42.5 57.5 72.5 10.0 15.0 15.0 20.0 17.5
GRPO 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 $0$ 0.0 2.5 2.5 5.0 5.0 5.0 2.5 2.5 5.0 5.0 5.0
[…]
MATH SFT 0.0 3.3 10.0 20.0 26.7 0.0 0.0 3.3 6.7 16.7 10.0 45.0 60.0 65.0 80.0 12.5 15.0 12.5 15.0 17.5
GRPO 0.0 13.3 16.7 20.0 36.7 0.0 0.0 0.0 3.3 6.7 20.0 37.5 40.0 52.5 70.0 25.0 30.0 30.0 30.0 32.5
[…]
QWEN
GSM8K SFT 0.0 0.0 0.0 0.0 6.7 0.0 0.0 0.0 6.7 6.7 0.0 10.0 17.5 52.5 62.5 0.0 2.5 7.5 7.5 7.5
GRPO 0.0 3.3 6.7 16.7 20.0 0.0 3.3 6.7 10.0 20.0 7.5 40.0 52.5 67.5 77.5 17.5 12.5 25.0 30.0 42.5
[…]
MATH SFT 0.0 0.0 0.0 0.0 6.7 0.0 0.0 0.0 0.0 3.3 0.0 12.5 35.0 47.5 65.0 0.0 2.5 7.5 7.5 7.5
GRPO 0.0 3.3 3.3 6.7 16.7 0.0 3.3 10.0 13.3 16.7 30.0 67.5 70.0 77.5 87.5 27.5 42.5 42.5 45.0 42.5
[…]
Table 1: Qwen and LLaMA 8B out-of-distribution evaluation using the best of GTPO, GRPO, and SFT on AIME 2024/25 and AMC, reporting pass@k (and maj@k for AMC), with $1\leq k\leq 64$.
@@ 6.1 Performance Evaluation
Training Dynamics of GTPO and GRPO. In Figure 2-top (the first two rows), we compare the training stability and performance of GTPO against GRPO. Formatting and accuracy rewards are reported as percentages, where the maximum value (100%) corresponds to a reward of 10 in both. On the GSM8K dataset with LLaMA, GTPO consistently outperforms GRPO across all training steps in both accuracy and formatting metrics. On the more challenging MATH dataset, GRPO (with $G=12$ and $\beta=0$) initially achieves slightly better accuracy around the midpoint of training. However, its performance drops sharply in the second half of training due to policy collapse, affecting both accuracy and formatting. In contrast, GTPO continues to improve steadily throughout training, avoiding collapse and maintaining stable performance. For Qwen 2.5, on both GSM8K and MATH, GTPO achieves comparable or improved accuracy relative to GRPO, with only a slight decrease in formatting performance (still above 97% in all runs). Note that, consistent with the analysis in Section 4.2, GRPO training curves for Qwen 2.5 are not affected by a policy collapse, as it shows high-entropy behavior. Overall, GTPO has more stable and reliable training compared to GRPO (see Fig. 5 in Appendix 9 for more details).
@@
In-distribution Evaluation. The trained LLMs were evaluated on the test sets of GSM8K and MATH using the pass@k pass@k and maj@k maj@k metrics. The former measures whether at least one of the top-$k$ completions yields a correct answer, while the latter assesses correctness via majority voting over the top-$k$ completions. Note that, to ensure a fair comparison, we evaluated GRPO on LLaMA using the model checkpoint corresponding to the training step with the highest accuracy reward, rather than the one obtained after policy collapse. Few curves are omitted because, even at their best checkpoint, the corresponding models were too unreliable and achieved consistently low performance. Figure 2 bottom (the last two rows) shows that GTPO consistently outperforms GRPO in almost all settings for both pass@k and maj@k, as $k$ varies from 1 to 32. This indicates that models trained with GTPO exhibit stronger self-consistency when answering questions (higher maj@k) and better coverage of the correct answer across multiple completions (higher pass@k). We also include testing comparisons with SFT, where GTPO consistently outperforms SFT in maj@k across all models and datasets, and achieves higher average performance in pass@k. Specifically, SFT surpasses GTPO only in pass@k for $k>5$ on MATH with LLaMA, but not in terms of correctness with maj@k. Interestingly, in GTPO, larger values of $G$ always lead to better performance, which is not the case for GRPO; in most cases, the latter benefits from a value of $\beta=10^{-6}$ compared to $\beta=0$.
Out-of-distribution Evaluation. We further evaluate the models on three out-of-distribution benchmarks: AIME 2024, AIME 2025, and AMC 2023. We report pass@k for $k\in\{1,8,16,32,64\}$, and for AMC also maj@k, replacing $k=1$ with $k=4$ since maj@1 is identical to pass@1. For each dataset, we select the GRPO and GTPO variants with the value of $G$ that achieved the highest pass@k on the in-distribution tests.
Overall, GTPO achieves the strongest generalization performance across both model families. For LLaMA, GTPO consistently outperforms SFT and matches or surpasses GRPO for almost all values of $k$ on AIME 2024, AIME 2025, and AMC 2023. The only exception is AIME 2025 at $k=64$ when trained on MATH, where SFT is slightly higher. For Qwen, GTPO obtains the best results on nearly all metrics and especially excels on AMC, where it achieves the highest pass@k and maj@k. GRPO occasionally approaches or slightly exceeds GTPO at very large $k$ (e.g., AMC pass@64 with MATH training), but SFT remains the weakest method overall, showing competitive performance only at large $k$.
@@
Figure 3: Accuracy and formatting of LLaMA on GSM8K. GRPO curves use different KL-$\beta$ values, while “No $\langle H\rangle_{i}$ – No $\delta_{i}$ for GTPO applies only the conflict-aware gradient correction.
(a) LLAMA GTPO train on MATH with different $\gamma$: entropy, accuracy, and formatting. “NO $\delta_{i}$” disables entropy filtering.
(b) Entropy, accuracy, and formatting in LLaMA trained with GTPO on GSM8K with different values of $\langle H\rangle_{i}$.
(c) QWEN on GSM8K: GRPO/GTPO with different $\langle H\rangle_{i}$ thresholds. OnlyNeg means $\langle H\rangle_{i}$ only for completions with $\mathcal{A}^{-}$.
(d) LLaMA trained on GSM8K with GRPO, and with GRPO/GTPO under the constraint $\langle H\rangle_{i}<1.05$ and $\delta_{i}=0.1$.
Figure 4: Comparison of entropy, accuracy, formatting under different training settings for GTPO and GRPO.